# CodeAssistBench (CAB): Dataset & Benchmarking for Multi-turn Chat-Based Code Assistance

### v2

@@ 4.2 Model Evaluation Results
Table 2: Correctness and Average Turns for each model, based on both original GitHub threads and CAB-generated conversations. Rec refers to recent repositories, while All includes all-time repositories.
Model Correctness (%) Avg. GitHub Turn (Human) Avg. CAB Turn
Correct Partially Corr. Incorrect Correct Partially Corr. Incorrect Correct Partially Corr. Incorrect
Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All
ChatGPT 4.1 Mini 16.49 29.14 30.41 36.86 53.09 34.00 1.72 2.77 2.17 3.56 2.18 3.05 2.94 2.35 3.66 3.33 5.70 4.28
DeepSeek R1 11.34 27.14 33.51 40.29 55.15 32.57 2.05 2.66 2.32 3.62 1.98 2.99 2.82 2.24 3.18 2.72 4.50 4.28
Llama 3.3 70B 9.33 13.58 25.91 40.75 64.77 45.66 1.67 2.72 2.02 3.48 2.21 3.02 3.22 2.68 4.06 3.49 4.67 4.50
Haiku 3.5 7.22 16.86 30.93 43.43 61.86 39.71 1.79 2.63 1.93 3.47 2.23 3.04 3.86 2.73 3.97 3.81 6.76 5.63
Sonnet 3.7 11.34 25.71 30.93 35.43 57.73 38.86 1.68 2.86 1.95 3.45 2.27 3.09 2.36 2.30 3.57 3.15 5.71 4.21
Sonnet 3.7 (Think) 13.40 27.43 27.32 38.86 59.28 33.71 1.81 2.50 2.17 3.76 2.14 3.00 2.50 2.20 3.25 2.85 4.95 4.26
We benchmarked six state-of-the-art large language models (LLMs) - ChatGPT 4.1 Mini, DeepSeek R1, Llama 3.3 70B, Haiku 3.5, Sonnet 3.7, and Sonnet 3.7 (Think mode) - as maintainer agents using the CAB dataset. Each model engaged in multi-turn conversations to resolve user issues, guided by satisfaction conditions and evaluated by an automated assessor.
4.2.1 Metrics
We evaluate model performance using two primary metrics: Correctness and Verbosity. For Correctness, we classify responses as: (1) Correct, successfully addresses all satisfaction conditions with a working solution; (2) Partially Correct, addresses some conditions or requires minor modifications; or (3) Incorrect, fails to address the core issue or provides non-working solutions. For Verbosity, responses are classified as: (1) Appropriate, provides necessary information concisely; (2) Verbose, contains unnecessary details or repetitions; or (3) Terse, lacks sufficient explanation. These classifications are determined by Sonnet3.7 acting as an automated judge agent, which evaluates satisfaction condition fulfillment, execution results (when applicable), and compares with reference solutions (see Appendix A.5 for the prompt).
@@ 4.2.2 Sampling Strategy
Given the multi-turn nature of our evaluation framework, running all 3,286 issues through interactive simulation would be computationally prohibitive. Furthermore, issue counts vary significantly across languages and turn lengths, which could bias the evaluation outcomes. To address these concerns, we constructed a balanced, representative evaluation subset using a two-stage sampling strategy. For each language, we first selected up to five issues containing Dockerfiles (where available), and then sampled additional issues to ensure ten samples per turn-length: 1,2,3,4, and 5+. When specific ranges had fewer than ten available issues, the remaining quota was reallocated to the preceding bucket. Finally, we obtained 350 datasets for the all-time repositories (50 per language) and 194 datasets for the recent repositories (C: 18; C#: 25; C++: 28; Java: 8; JavaScript: 39; Python: 50; TypeScript: 26).
4.2.3 Quantitative Metrics - Correctness
@@ 4.2.6 Language-Specific Analysis
4.3 Human Alignment Evaluation
Each GitHub issue was independently reviewed by three annotators. A total of 21 professional contractors participated in the evaluation, with an average of 5.1 years of programming experience. Annotators brought diverse technical backgrounds, spanning Python, JavaScript/TypeScript, C#, and Java, as well as familiarity with frameworks such as React, Django, Spring Boot, and .NET Core.
Overall, across 663 extracted conditions, 86.3% were judged accurate (i.e., correctly capturing what the user wanted), but only 65.7% were deemed complete (i.e., covering all of the user’s expectations). This indicates high precision - when a condition is extracted, it is usually correct - but lower recall, as annotators frequently identified missing satisfaction criteria. For example, in C++, 97.9% of extractions were accurate, yet 73.3% of cases required additional conditions. This suggests our pipeline favors conservative, reliable extractions over exhaustive coverage.
To put the 86.3% accuracy figure in context, we note that the average pairwise agreement among human raters was 82.4%, with some pairs achieving perfect agreement ($\kappa=1.0$) and others showing lower consistency. From the result, we interpret it as a strong result, indicating that the extracted conditions are nearly as reliable as those agreed upon by independent human annotators. Detailed annotation guidelines - including definitions of correctness and completeness, abstraction principles, and example outputs - are provided in Appendix G. Full per-language results and agreement statistics are available in Appendix F.
@@ 5 Limitations
While CodeAssistBench (CAB) offers a realistic and scalable framework for evaluating multi-turn programming assistance, it has several limitations. These include conservative condition extraction, limited evaluation coverage, templated user behavior, language scope constraints, and lack of statistical testing. We provide a full discussion in Appendix B.
@@ 6 Conclusion and Future Work
Our study introduces CodeAssistBench (CAB), a fully automated benchmark for evaluating multi-turn programming assistance grounded in real-world developer interactions. Unlike prior benchmarks that emphasize single-turn or synthetic tasks, CAB simulates realistic conversations over full codebases using containerized environments and issue-specific satisfaction criteria. Through extensive experiments, we find that state-of-the-art language models struggle with complex, multi-turn dialogues - particularly on recent repositories, where even the best models achieve only 16.49% correctness. CAB’s automated pipeline enables scalable dataset expansion while supporting rigorous, reproducible evaluation across languages, project types, and time periods, offering a more faithful measure of model capabilities than existing benchmarks.
Several promising directions could extend CAB’s impact, including improving satisfaction condition extraction to balance precision with recall, expanding language coverage beyond the current seven languages, incorporating more sophisticated user simulation strategies to better capture nuanced developer interactions, and developing specialized metrics for different assistance categories. We hope CAB facilitates deeper insights into the limitations of current systems and helps guide the development of future AI assistants that better support real-world software engineering workflows. We further discuss potential societal benefits and risks of CAB in Appendix C.
@@ A.3 Automated Dockerfile Synthesis and Repair
To ensure that every benchmarked issue can be built and tested in a fully self-contained environment, we automatically (i) generate an initial set of Dockerfile candidates and (ii) iteratively refine any candidate that fails to build.
@@
For each GitHub issue we collect a rich context bundle—repository URL, commit SHA, truncated issue body, README, a structure summary of the repo, up to two GitHub Actions workflows, and (if available) a reference Dockerfile from the same project. We then prompt the LLM to produce 5 self-contained Dockerfiles, each of which must (1) install dependencies, (2) clone the repo at the specified commit, and (3) build the project without executing the user’s code. The system prompt fixes the LLM’s persona (“expert Docker engineer”), while the user prompt supplies the per-issue context. By requesting plain text only (no Markdown fences) we can pipe the response directly to docker build.
@@
Phase 2: Fault-Directed Repair.
@@
A.4 Satisfaction-Condition Extraction
To evaluate whether an assistant’s reply actually meets a developer’s needs, we first distill each GitHub issue thread into a small set of user-satisfaction conditions—explicit criteria that any acceptable answer must fulfill. The extraction procedure is entirely automated and comprises two stages.
Stage 1: LLM-based Extraction.
@@ 4. NEED-FOCUSED (problem, not implementation)
Stage 2: Post-hoc Verification.
@@ 3. APPROPRIATE VERBOSITY
Appendix B Extended Discussion of Limitations
@@
Appendix C Broader Impacts
@@
Appendix D Hardware Requirements and Runtime Environment
@@
Appendix E Supplementary Material
@@
Appendix F Human Alignment Details
We conducted a targeted evaluation to assess the accuracy and completeness of automatically extracted user satisfaction conditions. We randomly sampled 10 GitHub issues per language across seven programming languages and collected annotations from three independent human raters. Each annotator evaluated two aspects: Condition Accuracy and Condition Completeness. Across 210 records, annotators reviewed a total of 663 satisfaction conditions. The detailed per-language breakdown is shown in Table 4.
@@ Appendix F Human Alignment Details
Appendix G Human Annotation Instructions

### v3

@@ 4.2 Model Evaluation Results
We benchmark six state-of-the-art LLMs as maintainer agents: ChatGPT 4.1 Mini, DeepSeek R1, Llama 3.3 70B, Haiku 3.5, Sonnet 3.7, and Sonnet 3.7 (Think). Models interact with simulated users until all satisfaction conditions are met or the turn limit is reached.
4.2.1 Evaluation Metrics
We measure:
  • Correctness: Correct, Partially Correct, or Incorrect, determined by a judge LLM evaluating condition satisfaction, execution outcomes, and reference alignment.
  • Verbosity: Appropriate, Verbose, or Terse, capturing explanation style and conciseness.
Judge prompts and scoring criteria are in Appendix A.5.
@@ 4.2.2 Sampling Strategy
Running all 3,286 issues in multi-turn mode is computationally prohibitive. We therefore construct a balanced subset: per language, up to five Dockerfile issues plus ten issues per turn length bucket (1, 2, 3, 4, 5+). Underflow buckets reallocate to earlier ones. This yields 350 datasets for All-Time repositories and 194 datasets for Recent repositories.
Table 2: Correctness and average CAB turns. Rec. = Recent repositories; All = All-Time.
Model Correctness (%) Avg. CAB Turn
Correct Partially Corr. Incorrect Correct Partially Corr. Incorrect
Rec. All Rec. All Rec. All Rec. All Rec. All Rec. All
ChatGPT 4.1 Mini 16.49 29.14 30.41 36.86 53.09 34.00 2.94 2.35 3.66 3.33 5.70 4.28
DeepSeek R1 11.34 27.14 33.51 40.29 55.15 32.57 2.82 2.24 3.18 2.72 4.50 4.28
Llama 3.3 70B 9.33 13.58 25.91 40.75 64.77 45.66 3.22 2.68 4.06 3.49 4.67 4.50
Haiku 3.5 7.22 16.86 30.93 43.43 61.86 39.71 3.86 2.73 3.97 3.81 6.76 5.63
Sonnet 3.7 11.34 25.71 30.93 35.43 57.73 38.86 2.36 2.30 3.57 3.15 5.71 4.21
Sonnet 3.7 (Think) 13.40 27.43 27.32 38.86 59.28 33.71 2.50 2.20 3.25 2.85 4.95 4.26
4.2.3 Correctness Results
@@
To better understand the reason behind this recency gap, we investigate two possibilities: (1) model knowledge limitations, and (2) intrinsic properties of newer repositories. To isolate these effects, we construct a lightweight synthetic ablation using 50 automatically generated repositories (details in Appendix J). Surprisingly, Claude 3.7 Sonnet achieved 74% correctness on these synthetic repositories compared to only 11.34% on recent real repositories. The synthetic repositories contained richer documentation and version constraints within model training windows, suggesting that the recency gap is primarily driven by post–knowledge-cutoff framework changes rather than properties of AI-generated code. Full methodology and analysis appear in Appendix J.
@@ 4.2.5 Conversation-Length Analysis
4.2.6 Verbosity Analysis
Beyond correctness, we assess whether models provide appropriately concise explanations. Figure 7 (Appendix) summarizes verbosity classifications across all models. Overall, models tend to err on the side of over-explanation: verbose responses account for 40–60% of outputs, while terse responses are rare (<5%). Sonnet 3.7 (Think) shows the most balanced behavior, producing the highest proportion of appropriately concise answers, while Haiku 3.5 and Llama 3.3 70B are the most consistently verbose. Interestingly, verbosity increases on recent repositories, likely due to higher uncertainty and evolving APIs. These results suggest that current LLMs prioritize thoroughness over brevity, and that verbosity patterns meaningfully vary across model families.
@@ 4.2.7 Language-Specific Analysis
4.3 Human Evaluation Studies
We conducted two human evaluation studies to validate CAB’s automated components: judge reliability and satisfaction condition quality.
4.3.1 Judge Validation
To assess the reliability of our automated LLM judge, two software engineers (3+ years experience) independently evaluated 310 model responses sampled across languages, difficulty levels, and issue types. The annotators achieved substantial inter-rater agreement (78.28%, Cohen’s $\kappa=0.68$; see Table 4 in Appendix B), establishing a strong human baseline.
To evaluate judge robustness across model upgrades, we additionally ran three successive versions of Anthropic’s Sonnet family—Sonnet 3.7, Sonnet 4.0, and Sonnet 4.5—as maintainer agents on the 149 issues where human annotators reached full agreement. Individually, Sonnet 3.7 produced 12 correct, 45 partially correct, and 87 incorrect responses; Sonnet 4.0 produced 18 correct, 38 partially correct, and 92 incorrect responses; and Sonnet 4.5 produced 18 correct, 48 partially correct, and 82 incorrect responses. The consistency of these distributions across model generations indicates that judge behavior is stable with respect to model choice, and that the verified subset is representative for downstream analyses.
Comparing the automated judge to human annotators, the LLM judge achieved 65.92% average agreement—84.2% of the inter-human baseline. Agreement was highest on objective dimensions such as verbosity (93.2% of human-level agreement) and lower on subjective dimensions such as technical correctness (86.7% of human-level agreement). Based on these results, we release the 149-issue verified subset (84.7% of the validation set) as a high-quality set of gold-standard labels for future research. Full methodology, annotation protocols, and judge–human comparison details are provided in Appendix B.
4.3.2 Error Category Analysis
To better characterize how models fail across different problem types, we categorized the 310 cases from our human annotation study (Section B) into seven coarse error types using LLM-assisted classification with Sonnet 4.5. Logic errors (28.8%) and environment configuration issues (28.5%) dominate the dataset, while performance-related issues exhibit the highest annotator–judge agreement (75.0%). The full error taxonomy, per-model breakdowns, and implications for future benchmark design are provided in Appendix K.
4.3.3 Satisfaction Condition Validation
We evaluated automatically extracted satisfaction conditions through annotations by 21 professional contractors (mean 5.1 years programming experience). Across 663 conditions from 70 randomly sampled issues, 86.3% were judged accurate but only 65.7% complete, indicating high precision with conservative recall. Our pipeline favors reliable extractions over exhaustive coverage. Detailed annotation guidelines and per-language statistics are in Appendices I and H.
@@ 5 Limitations
While CodeAssistBench (CAB) offers a realistic and scalable framework for evaluating multi-turn programming assistance, it has several limitations. These include conservative condition extraction, limited evaluation coverage, templated user behavior, language scope constraints, and lack of statistical testing. We provide a full discussion in Appendix C.
@@ 6 Conclusion and Future Work
Our study introduces CodeAssistBench (CAB), a fully automated benchmark for evaluating multi-turn programming assistance grounded in real-world developer interactions. Unlike prior benchmarks that emphasize single-turn or synthetic tasks, CAB simulates realistic conversations over full codebases using containerized environments and issue-specific satisfaction criteria. Through extensive experiments, we find that state-of-the-art language models struggle with complex, multi-turn dialogues - particularly on recent repositories, where even the best models achieve only 12.08% correctness on the verified subset. CAB’s automated pipeline enables scalable dataset expansion while supporting rigorous, reproducible evaluation across languages, project types, and time periods, offering a more faithful measure of model capabilities than existing benchmarks.
Several promising directions could extend CAB’s impact: improving satisfaction-condition extraction to balance precision and recall, expanding language coverage beyond the current seven languages, incorporating more sophisticated user-simulation strategies, and designing specialized metrics for assistance categories (Appendix L). We hope CAB enables deeper insights into current system limitations and guides the development of future AI assistants that more effectively support real-world software engineering workflows. We further discuss potential societal benefits and risks in Appendix D.
Acknowledgments
We thank Karthik N. R. from AWS AI Labs for improving the maintainability and usability of our codebase, and the broader AWS AI team for helpful discussions that shaped this work.
@@ A.3 Automated Dockerfile Synthesis and Repair
To ensure that every benchmarked issue can be built and tested in a fully self‑contained environment, we automatically (i) generate an initial set of Dockerfile candidates and (ii) iteratively refine any candidate that fails to build.
@@
For each GitHub issue we collect a rich context bundle—repository URL, commit SHA, truncated issue body, README, a structure summary of the repo, up to two GitHub Actions workflows, and (if available) a reference Dockerfile from the same project. We then prompt the LLM to produce 5 self‑contained Dockerfiles, each of which must (1) install dependencies, (2) clone the repo at the specified commit, and (3) build the project without executing the user’s code. The system prompt fixes the LLM’s persona (“expert Docker engineer”), while the user prompt supplies the per‑issue context. By requesting plain text only (no Markdown fences) we can pipe the response directly to docker build.
@@
Phase 2: Fault‑Directed Repair.
@@
A.4 Satisfaction‑Condition Extraction
To evaluate whether an assistant’s reply actually meets a developer’s needs, we first distill each GitHub issue thread into a small set of user‑satisfaction conditions—explicit criteria that any acceptable answer must fulfill. The extraction procedure is entirely automated and comprises two stages.
Stage 1: LLM‑based Extraction.
@@ 4. NEED-FOCUSED (problem, not implementation)
Stage 2: Post‑hoc Verification.
@@ 3. APPROPRIATE VERBOSITY
Appendix B Expert Validation: Detailed Methodology and Results
B.1 Validation Methodology
Dataset Construction.
We initially sampled 200 unique GitHub issues from real-world repositories across seven programming languages for expert validation. To ensure annotation quality, we filtered out 24 issues containing non-English content, as all annotators were English speakers, resulting in a final set of 176 unique issues.
We then generated AI responses using two state-of-the-art models for these 176 issues: GPT-4.1-mini and Claude Sonnet 3.5. Initially, both models generated responses for most issues (134 issues received responses from both models to enable comparative analysis). However, not all AI-generated responses were suitable for human annotation. We excluded responses where the AI output contained unicode character encoding issues or null fields in critical parts, which would have prevented annotators from accurately evaluating the responses. Specifically, 40 GPT responses and 38 Sonnet responses were excluded due to these data quality issues.
After this filtering, we retained 310 usable AI responses for human evaluation: 154 from GPT-4.1-mini and 156 from Claude Sonnet 3.5. Each of these 310 responses was independently evaluated by two expert annotators, yielding 620 total annotation records. The final language-wise distribution of validated AI responses per model is shown in Table 3.
Table 3: Language distribution of validated AI responses
Language GPT-4.1-mini Claude Sonnet 3.5
Python 41 42
JavaScript 30 31
C++ 25 25
C# 22 21
TypeScript 17 18
C 13 13
Java 6 6
Total 154 156
Annotation Protocol.
Two senior software engineers (Human1 and Human2) with 3+ years of experience each independently evaluated all 310 AI responses. Critically, annotators used the exact same evaluation prompts and criteria as the LLM judges to ensure fair comparison and minimize methodological confounds. Each response was evaluated across three dimensions:
  • Technical Correctness: Does the response accurately address the user’s technical question?
  • Verbosity: Is the response appropriately concise, verbose, or terse?
  • Overall Verdict: Is the response correct, partially correct, or incorrect?
Annotators worked independently without communication, and each AI response was evaluated by both annotators, yielding 620 annotation records (310 responses $\times$ 2 annotators).
B.2 Inter-Rater Reliability
The two expert annotators demonstrated substantial overall agreement at 78.28% (728/930 judgments across all dimensions; Cohen’s $\kappa=0.68$), establishing a robust baseline for evaluating LLM judge performance. According to established guidelines landis1977measurement, $\kappa$ values between 0.61–0.80 indicate substantial agreement, confirming that our annotation protocol achieves a high level of consistency. Agreement varied across evaluation dimensions, as shown in Table 4.
Table 4: Inter-rater reliability between Human1 and Human2
Dimension Agreement Percentage Cohen’s $\kappa$
Verbosity 263/310 84.84% 0.77
Technical Correctness 233/310 75.16% 0.63
Verdict 232/310 74.84% 0.62
Overall 728/930 78.28% 0.68
This variation reveals that while response style (verbosity) is relatively objective and unambiguous ($\kappa=0.77$, substantial agreement), assessing technical accuracy requires nuanced domain expertise and can yield different but equally defensible judgments ($\kappa=0.62\text{--}0.63$, moderate to substantial agreement). According to established guidelines landis1977measurement, $\kappa$ values between 0.61–0.80 indicate substantial agreement, and our results fall within this range across all dimensions. This level of agreement demonstrates that even experienced practitioners may weigh correctness, completeness, and stylistic factors differently when evaluating subjective technical tasks.
Our 78.28% agreement rate and Cohen’s $\kappa$ of 0.68 represent substantial inter-rater reliability, validating both our annotation protocol and confirming the inherent subjectivity in evaluating code quality and assistance effectiveness.
B.3 LLM-Expert Agreement
We compared automated LLM judge evaluations against both expert annotators to assess the reliability of LLM-as-judge for code assistance evaluation. Table 5 summarizes the agreement rates.
Table 5: Agreement rates between LLM judges and expert annotators
Judge Pair Overall Technical Verbosity Verdict
Human1 vs Human2 78.28% 75.16% 84.84% 74.84%
LLM vs Human1 70.22% 65.16% 79.03% 66.45%
LLM vs Human2 61.61% 53.87% 74.84% 56.13%
The LLM judges achieved 70.22% agreement with Human1, representing 89.7% of the inter-human baseline (78.28%), and 61.61% with Human2 (78.7% of baseline). The average LLM-human agreement of 65.92% reaches 84.2% of the inter-human agreement level, demonstrating that LLM judges approach human-level evaluation capabilities while maintaining scalability advantages. These results reveal several important findings:
Performance on Objective vs. Subjective Metrics.
LLM judges achieve their highest agreement with human experts on verbosity assessment (74–79%), where evaluation criteria are most clearly defined and objective. Agreement decreases for technical correctness (54–65%), where deep domain expertise and contextual understanding become critical. The 79.03% LLM-Human1 agreement on verbosity (93.2% of the 84.84% human baseline) demonstrates that automated evaluation excels at surface-level characteristics but requires improvement for nuanced technical assessment.
Variability in Human Standards.
The substantial 8.61 percentage point difference in LLM agreement between Human1 (70.22%) and Human2 (61.61%) indicates that individual annotator strictness, expertise emphasis, or judgment criteria significantly impact evaluation outcomes. This variability—comparable to the 8.06 point gap between Human1-LLM (70.22%) and Human1-Human2 (78.28%)—highlights the importance of multi-annotator validation and suggests that single-annotator studies may yield inconsistent conclusions. Our dual-annotator approach provides more reliable ground truth than typical single-expert evaluation.
B.4 Verified Dataset Release
Based on this validation, we release a verified subset of 149 issues (84.7% of the validation set) where at least one model response achieved full annotator concordance across all three evaluation dimensions. This verified subset provides high-quality ground truth for code assistance evaluation while maintaining broad coverage across programming languages and issue types: JavaScript (37), Python (37), TypeScript (23), C# (19), C++ (14), C (13), and Java (6). The strong verification rate across all languages validates the quality and clarity of our benchmark, providing researchers with reliable ground truth for benchmarking code assistance systems across diverse programming contexts. We plan to extend this verified subset in future work through additional expert validation.
Appendix C Extended Discussion of Limitations
@@
Appendix D Broader Impacts
@@
Appendix E Hardware Requirements and Runtime Environment
@@
Appendix F Docker Build Details
Most build failures in All-Time repositories stemmed from outdated dependencies or unavailable components, not from limitations of the LLM pipeline itself. To address this, CAB inferred project-specific requirements—such as relevant library versions, operating systems, and toolchains—by analyzing both the issue’s content and its creation timestamp, then tailored environment reconstruction accordingly via custom Dockerfile generation. The higher success rate in Recent repositories (97.6% vs. 78.2%) reflects better availability of current dependencies and more standardized build configurations in modern projects.
Appendix G Supplementary Material
@@
Appendix H Satisfaction Condition Validation Details
We conducted a targeted evaluation to assess the accuracy and completeness of automatically extracted user satisfaction conditions. We randomly sampled 10 GitHub issues per language across seven programming languages and collected annotations from three independent human raters. Each annotator evaluated two aspects: Condition Accuracy and Condition Completeness. Across 210 records, annotators reviewed a total of 663 satisfaction conditions. The detailed per‐language breakdown is shown in Table 7.
@@ Appendix H Satisfaction Condition Validation Details
Appendix I Human Annotation Instructions